OpenStudy (anonymous):

what is the standered form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0) ?

OpenStudy (princeharryyy):

(x-2)^2 + (y+3)^2 = r^2 plug in x = 2 and y = 0 in the equation above find r and finally replace the value of r in above equation with r

OpenStudy (princeharryyy):

@Jade3115 I guess you can solve it now. Else, ask again.

OpenStudy (princeharryyy):

r = 25 (x-2)^2 + (y+3)^2 = 25 answer